# Learning 3D-Gaussian Simulators from RGB Videos

### v1

@@ Abstract
Learning physics simulations from video data requires maintaining spatial and temporal consistency, a challenge often addressed with strong inductive biases or ground-truth 3D information – limiting scalability and generalization. We introduce 3DGSim, a 3D physics simulator that learns object dynamics end-to-end from multi-view RGB videos. It encodes images into a 3D Gaussian particle representation, propagates dynamics via a transformer, and renders frames using 3D Gaussian splatting. By jointly training inverse rendering with a dynamics transformer using a temporal encoding and merging layer, 3DGSim embeds physical properties into point-wise latent vectors without enforcing explicit connectivity constraints. This enables the model to capture diverse physical behaviors, from rigid to elastic and cloth-like interactions, along with realistic lighting effects that also generalize to unseen multi-body interactions and novel scene edits.
Figure 1: 3DGSim is a particle-based simulator that predicts scene dynamics from multi-view RGB images, leveraging recent advancements in point cloud transformers and 3D Gaussian scene representations. The predictions above showcase the model’s ability to capture deformations and shadows despite being trained on just 1,000 trajectories of six objects.
**footnotetext: Equal advising.
@@ 1 Introduction
Video generation models [1, 2, 3] hold great promise for enabling robotic reasoning in unstructured environments. However, despite advances in visual realism, their predictions can sometimes lack physical plausibility [4], exhibiting artifacts such as erratic object motion, unexpected shape changes, or lack of object permanence. These limitations reduce their reliability significantly, presenting a critical challenge for using these models in robot decision-making.
Table 1: Overview on recently proposed particle-based simulators. While most works resort to a combination of kNN and GNNs, our work distinguishes itself from prior work by resorting to 3D-GS, space filling curves for point cloud serialization, and training the inverse rendering encoder alongside a dynamics transformer.
Scene
representation
Inverse renderer
@@
Graph
synthesis
Dynamics
model
Forward
rendering
SDF-Sim [5] Mesh n.a. SDF GNN n.a.
Saleh et al. [6] Mesh n.a. Mesh GNN + Attention n.a
@@
FIGNet [8] Mesh faces n.a. BVH GNN n.a.
Robocraft [9] Point clouds n.a. (RGB-D) kNN GNN NeRF
VPD [10] Point clouds n.a. (RGB-D + UNet) kNN GNN NeRF
HD-VPD [11] Point clouds n.a. (RGB-D + UNet) kNN GNN + Transformer NeRF
@@
3DGSim (Ours) Gaussian splats MVSplat SFC Transformer 3D-GS
A strategy to enhancing a model’s physical accuracy is to embed physics-based inductive bias into its design. One realization of this idea involves utilizing a purely physics-based simulation combined with a learned mapping from images or text to the simulator’s state as in [13, 12, 14, 15]. While physics simulators [16, 17, 18, 19, 20, 21, 22] are an indispensable tool for robotics, fully leveraging their potential requires precise state estimation and meticulous dynamics tuning. Moreover, as noted by [4], learning a simulator from vision necessitates capturing complex physical phenomena, including material mechanics, light transport, magnetism, fluid dynamics, and thermodynamics. In turn, each bias introduced into vision-based simulators risks becoming a constraint, potentially limiting the model’s ability to generalize across scenarios. This aligns with Rich Sutton’s conclusion in his “bitter lesson”: over decades of AI research, models emphasizing scalability and adaptability have consistently outperformed those relying on handcrafted inductive biases and human-designed reasoning.
A promising direction to improve the physical accuracy of vision-based simulators are particle-based simulators (PBS) as shown in Table 1. Significant progress in language modeling has relied on identifying effective representation through learned token embeddings. Analogously, PBSs propose particles instead of pixels as a more suitable embedding for modeling dynamics. This approach is motivated by physical environments being inherently composed of atoms, whose interactions govern the appearance of a scene.
However, existing PBS structure point clouds using the k-nearest neighbors (kNN) algorithm and propagate information through graph neural networks (GNNs), introducing scalability limitations [11] and inductive biases that hinder learning. To address these challenges, we propose 3DGSim, a novel PBS framework that jointly trains inverse rendering and dynamics model directly from RGB images using 3D Gaussian splatting. Specifically, 3DGSim introduces the following key contributions:
  • Training on RGB video: Avoiding access to meshes or depth cameras, 3DGSim learns particle interactions directly from multi-view RGB images by representing scenes as 3D Gaussian point clouds.
  • Point cloud serialization over kNN: Instead of relying on kNN to identify spatially close points, 3DGSim leverages temporal point cloud serialization, significantly enhancing model scalability.
  • Temporal encoding and merging layer: Replacing GNN-based message passing, 3DGSim utilizes temporal encoding and merging of serialized point clouds, allowing transformers to process unstructured point-cloud embeddings efficiently.
  • Inverse physics rendering: 3DGSim jointly trains inverse rendering alongside the dynamics model, embedding motion priors directly into the 3D Gaussian representation while mapping these latent embeddings to splats only before rendering.
  • Open-source datasets and code.
2 Related work
Given the vast scope of data-driven simulators, we focus on vision-based simulations of the real world using particle-based representations. These particle-based simulators (PBS), as summarized in Table 1, vary in how they generate scene representations and in their approaches to propagating particles over time.
@@ 2 Related work
Common scene representations include point clouds (aka particles), meshes, signed distance functions (SDFs), neural radiance fields (NeRFs) [23], and 3D Gaussians (aka splats) [24]. Point clouds – being 3D points located near a scene’s objects – can be obtained directly from RGB-depth cameras as done in [9, 10, 11] or via inverse rendering. Notably, [10, 11] deploy U-Net [25] inspired encoder architectures to obtain pixel-wise features that are trained jointly with the dynamics model. As highlighted by [10], this joint training approach, in contrast to the independent training used in [1], enables the encoder to capture features that are relevant to the system’s dynamics. We consider this aspect crucial for the design of PBS, therefore also train the encoder alongside dynamics via MVSplat [26]. Many works [7, 9, 10, 11, 12] on PBS resort to NeRFs for rendering images, which is an implicit representation that stores rendering cues in the weights of the neural network. In comparison, we represent visual information directly in the particle clouds as 3D Gaussians (aka splats). Splats are an explicit representation that yields high-fidelity renderings while also being significantly faster compared to NeRF-based ray casting [24].
Using point connectivity for learning dynamics
While the physically grounded representation of PBS is its core strength, the unstructured nature of these point clouds in both spatial and temporal dimensions presents a major challenge. To connect points spatially, [8, 6] require object meshes to establish connectivity within objects. [8] resorts to the BVH [27] to connect faces between bodies. As tree search methods such as BVH are computationally expensive iterating over all vertices, [5] motivates learning an SDF for collision detection.
While meshes can be obtained from point clouds using meshing algorithms [28], such algorithms are computationally expensive and do not solve the problem of connecting points between objects. Therefore, most PBS [29, 7, 10, 11, 9, 12] resort to furthest-point-sampling (FPS) and k-nearest neighbors (kNN) for connecting neighboring particles and deploy graph neural networks (GNN) for predicting dynamics. However, kNN is expensive requiring a forward pass on $KN$ particle-neighbor pairs with number of neighbors $K$ and number of particles $N$. Furthermore, kNN connects particles only locally such that passing information through a graph requires many parsing steps. However, this approach is problematic for rigid bodies, for which instant changes in velocity require passing information through the whole graph in a single time step. Recent works [6, 11] suggests adding attention layers to efficiently pass information through the graph.
Also closely related is DEL [12], which uses the material point method as inductive bias. Naturally, this physics bias also limits its applicability for the many real-world physical phenomena that this model cannot capture. In addition, DEL requires small time steps for stability, incurs considerable per-step kNN computation costs, and relies on accurate velocity estimates.
In contrast to these approaches, we employ point cloud serialization using space-filling curves, effectively reducing the computational overhead associated with kNN. These serializations are fed to a sparse convolutional neural network to obtain patch-wise particle embeddings. Further, drawing inspiration from [6, 11], we employ PTv3 [30], a hierarchical UNet-like Point Transformer as dynamics model, avoiding physics biases. By performing patch-wise pooling followed by attention, we avoid using global graph nodes that could act as a bottleneck for passing particle-wise features.
Tracking
Tracking and dynamics models serve distinct purposes in point cloud processing. Tracking focuses on estimating the current scene representation using present (and optionally past) images, often by identifying and following keypoints. In contrast, dynamics models predict future representations based on past observations, capturing the underlying physics that govern the scene’s motion. For instance, [31, 32, 33, 34] employ Gaussian Splatting, primarily for keypoint tracking and modeling of static deformations as well as robot control [35, 36, 37].
3 3DGSim
As illustrated in Figure 2, 3DGSim consists of three building blocks: (i) an encoder that maps multi-view RGB images to 3D Gaussian particles, (ii) a dynamics model that simulates the motion of the Gaussians through time, and (iii) a renderer that yields images from the Gaussians. These components are trained jointly end-to-end to perform image reconstruction.
Figure 2: 3DGSim receives multi-view RGB images of the scene from $T$ preceding time steps and is trained to predict the next image. First, an inverse renderer maps each set of images to 3D Gaussian point clouds. The Gaussian point clouds are then transformed by the dynamics transformer containing a temporal merging layer to predict a future point cloud. Finally, 3D-GS renders images from the predicted Gaussian point clouds.
3.1 Forward rendering
@@
3DGSim uses a 3D Gaussian point cloud representation which has been chosen to enable fast image rendering. Rendering is done using parts of the 3D Gaussian Splatting framework [24] which has emerged as an effective method for multi-view 3D image reconstruction, offering rapid inference, high fidelity, and the ability to edit scenes. Gaussian splatting uses a set of 3D Gaussian primitives, each parameterized by
@@
with the Gaussian’s mean $p_{i}$ (aka particle position), its rotation $r_{i}$, spherical harmonics $c_{i}$ (defining coloring), scale $s_{i}$, and opacity $\sigma_{i}$. To render novel views, these primitives are projected onto a 2D image plane using differential tile-based rasterization. The color value at pixel $\mathbf{p}$ is calculated via alpha-blend rendering:
$I(\mathbf{p})=\sum_{i=1}^{N}\alpha_{i}c_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$
where $\alpha_{i}=\sigma_{i}e^{-\frac{1}{2}(\mathbf{p}-p_{i})^{{}^{\top}}\Sigma_{i}^{-1}(\mathbf{p}-p_{i})}$ is the 2D density, $I$ is the rendered image, $N$ is the number of primitives in the tile, and $\Sigma_{i}$ is the covariance matrix given by $\Sigma_{i}=r_{i}s_{i}r_{i}^{{}^{\top}}$ for improved computational stability.
3.2 Inverse rendering encoder
As an encoder, we employ MVSplat [26] to obtain a scene representation in form of 3D Gaussians. MVSplat deploys a feed-forward network $f_{\phi}$ with parameters $\phi$ that maps $M$ images =$\mathcal{I}=\{I^{m}\}_{i=m}^{M}$ with $I^{m}\in\mathbb{R}^{(H\times W\times 3)}$ to a set of 3D Gaussian primitives
@@ 3.2 Inverse rendering encoder
Figure 3: As encoder, we use MVSplat which estimates a 3D Gaussian splat for every image pixel in the first set of multi-view cameras. The Gaussians are estimated by projecting them onto the second set of cameras planes.
Latent particle-wise features
Insipred by [11], we extend MVsplat to produce point-wise latent features
$\tilde{g}_{i}=(p_{i},f_{i}),$ (2)
where the explicit rendering parameters typically associated with 3D Gaussian splatting $g_{i}$ are replaced or supplemented by pixel-wise feature vectors $f_{i}$. These latent features capture latent scene properties and can optionally enrich the explicit representation $g_{i}$.
We defer the generation of explicit rendering parameters to the end of the pipeline, before rasterization. Specifically, we introduce a pixel-wise MLP that decodes latent features $f_{i}$ into explicit 3DGS rendering parameters $g_{i}$. This allows the pixel-wise features $f_{i}$ to better capture physics-related per-particle properties and avoids constraining the encoder with predicting only rendering-specific parameters.
MVsplat initially projects image-space parameters into the 3D world coordinate frame using camera intrinsics and estimated depth candidates, thus establishing accurate spatial scaling. Consequently, naive approaches for directly generating latent vectors $\tilde{g}_{i}$ typically fail. Instead, we propose generating the pixel-wise latent features $f_{i}$ via a dedicated MLP conditioned on projected parameters, depth, pixel shift, densities and geometric context captured through the Plücker coordinates of viewing rays:
$\mathbf{r}=[\,\mathbf{o}\times\mathbf{d}\,\mid\,\mathbf{d}\,],$
where $\mathbf{o}$ and $\mathbf{d}$ are the ray’s origin and direction, respectively [38]. The geometric encoding $\mathbf{o}\times\mathbf{d}$ captures complex spatial relations otherwise challenging to parameterize. Furthermore, our latent features $f_{i}$ explicitly decompose into a dynamic component $f_{i}^{\text{dyn}}$, which the dynamics model can modify, and an invariant component $f_{i}^{\text{inv}}$, which remains fixed across all simulation steps. This decomposition provides flexibility, allowing the dynamics model to selectively update only $f_{i}^{\text{dyn}}$, while preserving $f_{i}^{\text{inv}}$. For clarity, even though we may subsequently refer simply to ”updating $f_{i}$” in the text, we always update exclusively the dynamic component $f_{i}^{\text{dyn}}$, leaving the invariant component $f_{i}^{\text{inv}}$ unchanged.
Masking background Gaussians
For each time step $t_{k}$, the encoder predicts for every image pixel a 3D Gaussian with latent feature $\tilde{g}_{i}$. In turn, for $M$ images of size $H\times W$ the total number of 3D Gaussians would amount to $M\times H\times W$. As not all Gaussian are equally important for predicting object motion as illustrated in Figure 3, Gaussians that belong to the scene’s background can be removed from the representation to save compute such that total number of Gaussian per point clouds amounts to $N_{k}$. Further, as done in [10], Gaussians that belong to static surfaces can be fed to the model but will not receive updates.
3.3 Dynamics transformer
The dynamics model predicts
$\Delta p_{T}=p_{T+1}-p_{T}\text{ and }\Delta f_{T}=f_{T+1}-f_{T},$ (3)

### v2

@@ Abstract
Realistic simulation is critical for applications ranging from robotics to animation. Learned simulators have emerged as a possibility to capture real world physics directly from video data, but very often require privileged information such as depth information, particle tracks and hand-engineered features to maintain spatial and temporal consistency. These strong inductive biases or ground truth 3D information help in domains where data is sparse but limit scalability and generalization in data rich regimes. To overcome the key limitations, we propose 3DGSim, a learned 3D simulator that directly learns physical interactions from multi-view RGB videos. 3DGSim unifies 3D scene reconstruction, particle dynamics prediction and video synthesis into an end-to-end trained framework. It adopts MVSplat to learn a latent particle-based representation of 3D scenes, a Point Transformer for particle dynamics, a Temporal Merging module for consistent temporal aggregation and Gaussian Splatting to produce novel view renderings. By jointly training inverse rendering and dynamics forecasting, 3DGSim embeds the physical properties into point-wise latent features. This enables the model to capture diverse physical behaviors, from rigid to elastic, cloth-like dynamics, and boundary conditions (e.g. fixed cloth corner), along with realistic lighting effects that also generalize to unseen multibody interactions and novel scene edits. For more information, visit: https://mikel-zhobro.github.io/3dgsim/.
@@ 1 Introduction
Figure 1: 3DGSim works directly on multi-view RGB videos and is trained end-to-end on next image prediction. The dynamics model (transformer) operates on particles with latent feature. A learned mapping transforms them into 3D Gaussian Splats for novel view rendering.
Simulating visually and physically realistic environments is a cornerstone for embodied intelligence. Robots must soon tackle tasks like opening washing machines, folding laundry, or tending plants. Traditional analytical simulators demand exact geometry, poses, and material parameters, making arbitrary scene simulation impractical. An alternative is to learn models that predict future states of a scene in large scale observations, as evidenced by the striking visual realism of 2D video generation methods (Li et al., 2022; Wu et al., 2023; NVIDIA et al., 2025). However, pure 2D approaches lack 3D structure awareness, leading to failures in occlusion handling, object permanence, and physical plausibility (Motamed et al., 2025).
3D-based representations address many of these shortcomings, as shown by recent learned simulators which model a wide range of physical phenomena, from fluids and soft materials to articulated and rigid body dynamics Li et al. (2019); Allen et al. (2023). Yet, scaling such methods to data-rich regimes remains challenging as most methods require privileged signals (object-level tracks, depth sensors, physics prior) or hand-crafted graph constructions.
To bridge this gap, we identify three pillars for generalizable, scalable visuo-physical simulation from videos: (1) 3D visuo-physical reconstruction from raw RGB observations; (2) Minimal-bias dynamics modeling to capture diverse physics without explicit biases; (3) Efficient, differentiable decoding back to image space for supervision via reconstruction loss from videos.
Graph neural networks (GNNs) (Sanchez-Gonzalez et al., 2020; Xue et al., 2023; Whitney et al., 2023, 2024; Shi et al., 2024; Wang et al., 2024) have shown great promise in introducing relational inductive biases to handle the unstructured nature of particle sets. This has allowed GNN-based particle simulators to make major progress on all three pillars. In particular, Whitney et al. (2023) jointly train an encoder and dynamics model to learn visuo-physical pixel features from RGBD, and in follow-up work (Whitney et al., 2024) eliminate point correspondences via abstract temporal nodes or per-step models with merging. Driess et al. (2023) demonstrate end-to-end dynamics training of composable NeRF fields from raw RGB images. These advances, in combination with recent advances in feed-forward inverse rendering Chen et al. (2024) and fast differentiable rendering of particles Kerbl et al. (2023), encourage us to ask the question: can we give up the inductive bias arising from locally connected graphs and still learn 3D particle-based simulators?
To this end, we build 3DGSim, a fully end-to-end differentiable framework that embraces the power of scalable computation over hand-crafted biases. 3DGSim begins by inferring 3D visuo-physical features from raw multi-view RGB images through a feed-forward inverse renderer based on MVSplat, extended with a feature extraction module. We then introduce a transformer-only dynamics engine, avoiding kNN-based graph construction and manually designed edge features in favor of learned spatiotemporal embeddings. Finally, a Gaussian Splatting head enables training on an image reconstruction loss from multi-view videos.
Specifically, 3DGSim introduces the following key contributions:
  • Inverse Renderer: Extends MVSplat with a feature extraction module fusing pixel-aligned features into a particle visuo-physical latent representation.
  • Temporal Encoding & Merging Layer: Discards abstract temporal nodes in favor of a hierarchical module that processes an arbitrary number of timesteps.
  • Transformer-Only Dynamics Engine: Removes graph biases and instead uses space-filling curves and learned embeddings for particle-based simulation.
  • End-to-End Differentiable Framework: Connects inverse rendering, transformer dynamics, and Gaussian splatting-based decoding to training for next-frame image reconstruction.
  • Open Source Release: We release the code and dataset to establish a reproducible baseline for future visuo-physical simulation research.
2 Related work
Table 1: Overview on recently proposed particle-based simulators. While most works resort to a combination of kNN and GNNs, our work distinguishes itself by resorting to 3D Gaussian Splatting, space filling curves (SFC) for point cloud serialization, and training the inverse rendering encoder alongside a dynamics transformer.
{NiceTabular}
@@ 2 Related work
[…]
3DGSim (Ours) Gaussian splats MVSplat SFC Transformer 3DGS
@@
Common scene representations include point clouds (particles), meshes, signed distance functions (SDFs), neural radiance fields (NeRFs) Mildenhall et al. (2021), and 3D Gaussians (splats) Kerbl et al. (2023). Point clouds, which approximate object surfaces, can be obtained from RGB-D sensors Shi et al. (2024); Whitney et al. (2023, 2024) or via inverse rendering Wang et al. (2025); Murai et al. (2024); Chen et al. (2024). Several works Whitney et al. (2023, 2024) use U-Net–style encoders trained jointly with the dynamics model, allowing the extracted features to be optimized for physical prediction, a strategy shown to outperform independently trained encoders Li et al. (2022). We adopt this joint training approach using MVSplat Chen et al. (2024), where the encoded features are initially bound to camera parameters. To unify these visuo-physical latents in a global frame, we introduce a learned feature transformation module that maps them into a consistent 3D representation. While many PBS methods render from NeRFs Xue et al. (2023); Shi et al. (2024); Whitney et al. (2023, 2024); Wang et al. (2024); Driess et al. (2023), we instead encode visual appearance directly in the particle cloud using 3D Gaussians. This explicit representation offers high rendering fidelity and significantly improved efficiency over NeRF-based rendering Kerbl et al. (2023), supporting scalability.
GNN based particle-based simulators (PBS)
Graph neural networks (GNNs) introduce relational inductive biases well-suited for modeling the unstructured nature of particle systems. Early work Sanchez-Gonzalez et al. (2020); Li et al. (2019) demonstrated that GNN-based PBS can fit trajectories across a range of physical phenomena. However, GNNs struggle with rigid bodies, where instantaneous velocity changes require long-range message passing across the entire graph in a single step. To address this, later works incorporate mesh structures Pfaff et al. (2021); Allen et al. (2022b) or signed distance functions (SDFs) Rubanova et al. (2024) to enforce object-level coherence. While effective in rigid-body settings, these methods do not generalize easily to deformable or fluid systems. Recent works Saleh et al. (2024); Whitney et al. (2024) suggest adding attention layers to efficiently pass information through the graph. Wang et al. (2024) move toward greater data efficiency by incorporating physics-inspired biases such as the Material Point Method, though limiting broad applicability and requiring small time deltas to simulate. To address temporal correspondence, Whitney et al. (2023) introduces abstract temporal nodes, while Whitney et al. (2024) combines GNNs with transformers to improve memory efficiency by processing and merging pairs of timesteps. However, the method is restricted to two-step horizons, as extending it requires training a separate model for each additional timestep. Typically, these GNN based methods rely on kNN within a fixed radius and hand-crafted features based on object associations and distances to define graph edges and edge features. Message passing and spatial pooling via furthest-point-sampling (FPS) are then used to aggregate information for dynamics prediction. However, kNN and distance computations are expensive and take up 54% of the forward time Wu et al. (2024a), which limits scalability and prevents real-time forecasting. In contrast, we follow the design of PTv3 Wu et al. (2024a). In 3DGSim, we trade off exact KNN neighborhood computation with space-filling curve–based ordering of particles and use sparse convolutions to encode relative positions, avoiding distance calculations. To scale temporally, we propose Temporal Merging with Grid Pooling to construct a hierarchical spatiotemporal, UNet-style Point Transformer for dynamics prediction, abandoning hand-crafted physics in favor of learned representations.
Analytical particle simulators as physical prior
Our work differs in purpose from applications which use Gaussian Splatting particles and analytical PBS as physical prior (e.g. of the shelf differentiable MPM simulator) to accomplish a series of tasks such as tracking Luiten et al. (2024); Keetha et al. (2024); Zhang et al. (2024a); Abou-Chakra et al. (2024), dynamic scene reconstruction Wu et al. (2024b); Huang et al. (2023); Yu et al. (2023) or animation Xie et al. (2023); Zhang et al. (2024b); Lin et al. (2025). While those models can be extended to do parameter estimation via the differentiable nature of simulators Abou-Chakra et al. (2024) they do not fit the generalization of our problem setting.
3 Preliminaries
3DGSim is build atop several prior works, namely: 3D-Gaussian splatting which enables fast rendering, MVSplat which yields 3D Gaussian point clouds from multi-view images, and PTv3 which enables efficient neural processing of 3D point clouds.
Gaussian Splatting
@@ 3 Preliminaries
3D Gaussian splatting (3DGS) Kerbl et al. (2023) is an effective framework for multi-view 3D image reconstruction, representation, and fast image rendering and gained rapid popularity due to its support for rapid inference, high fidelity, and editability of scenes. Gaussian splatting uses a collection of 3D Gaussian primitives, each parameterized by
@@
with the Gaussian’s mean $p_{i}$ (particle position), its rotation $r_{i}$, spherical harmonics $c_{i}$ (defines coloring), scale $s_{i}$, and opacity $\sigma_{i}$. To render novel views, these primitives are projected onto a 2D image plane using differential tile-based rasterization. The color value at pixel $\mathbf{p}$ is calculated via alpha-blend rendering: $I(\mathbf{p})=\sum_{i=1}^{N}\alpha_{i}c_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ where $\alpha_{i}=\sigma_{i}e^{-\frac{1}{2}(\mathbf{p}-p_{i})^{{}^{\top}}\Sigma_{i}^{-1}(\mathbf{p}-p_{i})}$ is the 2D density, $I$ is the rendered image, $N$ is the number of primitives in the image, and $\Sigma_{i}$ is the covariance matrix given by $\Sigma_{i}=r_{i}s_{i}r_{i}^{{}^{\top}}$ for improved computational stability.
Figure 2: MVSplat uses a cost volume with plane sweeping to regress pixel-wise 3D Gaussians, which are unprojected to world frame using camera parameters.
MVSplat: Multi-view images to 3D point clouds
MVSplat deploys a feed-forward network $f_{\phi}$ with parameters $\phi$ that maps $M$ images $\mathcal{I}=\{I^{m}\}_{i=m}^{M}$ with $I^{m}\in\mathbb{R}^{(H\times W\times 3)}$ to a set of 3D Gaussian primitives (Fig.˜2)
@@
4 3DGSim
3DGSim is a fully differentiable pipeline that, given $T$ past multi-view RGB frames, reconstructs 3D particles with latent features, simulates their motion, and renders the next frames. It consists of three jointly trained modules (Fig.1): (i) an encoder that maps multi-view RGB images to 3D particles, (ii) a dynamics model that simulates the motion of these particles through time, and (iii) a renderer that yields images by first mapping the particles to Gaussian splats.
4.1 State Representation
To simulate physical scenes from vision, we require a state representation that is both expressive enough to capture fine-grained 3D and physical properties, and compact enough to enable efficient learning and prediction. While an explicit 3DGS representation $g_{i}(t_{k})$ offers geometric and visual completeness, it is insufficient for dynamics modeling. Instead, we distill the state of each particle into a more compact representation:
$\tilde{g}_{i}(t_{k})=\bigl{(}p_{i}(t_{k}),\,f_{i}(t_{k})\bigr{)}$ (2)
where $t_{k}$ denotes the k-th timestep and $f_{i}\in\mathbb{R}^{d}$ the visuo-physical latent particle feature, encoding shape, appearance, and dynamic properties. Unless otherwise stated, we omit the timestep $t_{k}$ and the particle index $i$ when the statement applies to all timesteps or particles, respectively.
Optional: Masking and Freezing of Particles
At each timestep $t_{k}$, the encoder yields pixel-aligned features for each input image. As an optional step, we apply a foreground mask to discard particles likely belonging to the static background, retaining a reduced set of $N_{k}$ particles per time step (Fig. 2). Additionally, as originally suggested by Whitney et al. (2023), static particles can optionally be “frozen”, i.e. act as input to the dynamics model but are excluded from position updates. These optional strategies improve efficiency without being necessary for successful training, as shown in Sections˜5 and B.
Figure 3: Position $p$ and dynamic features $f^{\text{dyn}}$ are updated while $f^{\text{inv}}$ remain constant.
Invariant and dynamic feature decomposition
We decompose each particle’s visuo-physical feature into an invariant and a dynamic part as shown in Fig.˜3, writing
$f_{i}=f_{i}^{\text{inv}}\oplus f_{i}^{\text{dyn}},$
where $\oplus$ denotes concatenation. The dynamics model updates only $f_{i}^{\text{dyn}}$, while leaving $f_{i}^{\text{inv}}$ unchanged. For clarity, we will refer to the dynamics update as “updating $f_{i}$”, though only the dynamic component $f_{i}^{\text{dyn}}$ is altered.
4.2 View-Independent Inverse Renderer
In the proposed augmentation of MVSplat, the pixel-aligned features $\hat{f}_{i}^{\prime}$ are tied to the view-dependent camera frame in which they were observed. This camera-centric encoding hinders generalization and leads to subpar model performance as dynamics prediction should ideally be invariant to the observer’s viewpoint. To address this, 3DGSim extends MVSplat with a learned feature encoding network that transforms pixel-aligned features $\hat{f}_{i}^{\prime}$ into view-independent latent features $f_{i}$. The feature encoding network leverages the pixel depth, pixel shift, density, and ray geometry, represented via Plücker coordinates Plücker (1868-1869) to infer spatially consistent 3D features via FiLM conditioning Perez et al. (2017). This yields an inverse rendering encoder that outputs canonically anchored particle states suitable for downstream dynamics learning. Details are provided in Section˜A.1.
4.3 Dynamics Model
At the core of our method is the dynamics model, a transformer architecture operating on particle sets in space and time. The dynamics model receives as input the full particle trajectory
$\big{\{}\{\tilde{g}_{i}(t_{k})\}_{i=1}^{N_{k}}\big{\}}_{k=1}^{T},\quad\text{where }\tilde{g}_{i}(t_{k})=\big{(}p_{i}(t_{k}),f_{i}^{\text{inv}}(t_{k}),f_{i}^{\text{dyn}}(t_{k})\big{)},$ (3)
and predicts the updated dynamic features at the next timestep
$\Delta p(t_{T}),\Delta f^{\text{dyn}}(t_{T})=\mathrm{Dynamics\,Model}\left(\{\{\tilde{g}_{i}(t_{k})\}_{i=1}^{N_{k}}\}_{k=1}^{T}\right),$ (4)